# Systematic Reward Gap Optimization for Mitigating VLM Hallucinations

### v3

@@ Appendix B Curriculum Learning
Effectiveness of Hard Mining. To investigate the impact of curriculum learning on model’s ability to avoid subtle errors, we conduct a quantitative analysis comparing effectiveness in reducing hallucination across different types of varying difficulties. The RefoMB benchmark [18] is used for this ablation, as its detailed categorizations allow us to pinpoint the benefits of curriculum learning more precisely. The results in Figure 7 clearly demonstrate the progressive improvements achieved through our proposed TPR method and curriculum learning strategy. On challenging hallucinations for baseline LLaVA-1.5 model, i.e., “Quantities” (16.7%) and “Spatial Relations” (14.3%), the introduction of curriculum learning in TPR-CL yields further substantial improvements, providing an additional +20.8/+35.7 point increase over its greedy variant. For “Existence” and “Attributes”, where the greedy variant already performed strongly, TPR-CL still offers valuable refinements, with additional gains of +5.3 and +8.3 points respectively. These findings underscore the effectiveness of the hard mining component within our curriculum learning strategy. This is especially crucial for tackling more subtle and complex hallucinations related to quantities and spatial relationships, where the baseline model and even the greedy TPR approach show limitations.
@@
Appendix C Cost-Effectiveness
To quantify the cost-effectiveness of our proposed TPR, Figure 5 provides a comparative illustration against rewriting-based and ranking-based methods. For this analysis, we simulate rewriting-based approaches by using GPT-4V [2] as an external rewriter to generate 8,000 preference data instances for model alignment. To avoid introducing external hallucination patterns inconsistent with the base model’s intrinsic failure modes, we only correct hallucinations present in $y_{w}$, while original responses serve as $y_{l}$. The computational overhead for this rewriting-based method is determined from OpenAI API billing. The computational overhead for TPR and the ranking-based RLAIF-V [18], which typically utilize open-source models for data curation, is estimated based on API pricing from cloud platforms such as OpenRouter.
As Figure 5 demonstrates, our proposed TPR exhibits the most favorable trajectory, achieving a rapid and substantial reduction in hallucinations at minimal cost. While the GPT-4V rewriting approach also yields a significant reduction, it requires a generally higher cost. Notably, TPR can achieve a comparable level of hallucination reduction at roughly one-third of the cost required by the GPT-4V approach (approximately \$10 for TPR  vs. over \$30 for GPT-4V). In the other hand, the ranking-based RLAIF-V displays a much gentler decline in hallucination levels relative to cost. This suboptimal cost-effectiveness might stem from its reliance on ranking potentially flawed model outputs, which can lead to low informativeness or insufficient reward gaps to penalize hallucinations. And for ranking-based methods (RLAIF-V), its curve shows a much gentler decline in hallucination levels relative to cost, probably because using flawed model outputs for ranking leads to low informativeness or insufficient reward gaps to penalize the hallucinations. These findings clearly indicate that TPR not only achieves lower overall hallucination levels than both rewriting- and ranking-based methods but at a significantly lower computational cost, highlighting its practical advantages for efficient VLM alignment.
Appendix D More Ablation Studies
Self-Labeling: Using the Reference Model Itself as Labeler. The approach of scoring fine-grained semantic units within our intra-topic ranking mechanism potentially alleviates the demand for an exceptionally capable labeler model [18]. This prompted an exploration into the effectiveness of using the reference model itself for scoring and ranking after self-resampling, thereby investigating the boundaries of model self-improvement. Accordingly, we introduce TPR-SL (Self-Labeling), a variant of TPR that leverages the base LLaVA-1.5-7B model [7] as its own preference labeler, instead of a more powerful one like LLaVA-NeXT-34B [4] used in our main experiments. The performance of TPR-SL compared to the LLaVA-1.5-7B baseline is presented in Table 3. Specifically, TPR-SL achieves substantial improvements over the baseline LLaVA-1.5-7B across both hallucination mitigation and general VLM capabilities. These findings are significant as they highlight the effectiveness of TPR in a self-labeling scenario. It validates a practical and efficient paradigm for VLM self-improvement that minimizes reliance on human annotations or powerful external models.
Hallucination Benchmarks General Benchmarks
Model ObjHal MMHal AMBER RefoMB POPE LLaVA-B MMstar
@@ Appendix D More Ablation Studies
LLaVA-1.5-7B 53.6 25.2 2.36 51.0 73.5 77.6 30.8 12.1 85.9 59.7 30.3
+ TPR-SL 5.8 3.0 2.67 44.8 81.7 86.7 53.5 30.3 86.1 73.7 32.8
Qwen-VL-2B [5] 42.4 36.3 2.85 47.9 68.3 81.0 48.5 20.7 86.5 83.2 29.1
+ Naive RLAIF 36.5 31.7 2.81 49.0 70.9 81.8 52.5 23.7 87.0 83.2 30.3
+ TPR-SL-2B 19.5 13.7 2.98 43.8 74.3 84.3 56.6 28.3 87.1 83.8 31.2
Generalization on Different Model Architecture. To verify the generalizability of the TPR paradigm, we apply it to a different model architecture, Qwen-VL-2B [5]. Specifically, Qwen-VL-2B serves as the reference model for both generating topic-level alternatives and subsequent selective replacement, with intra-topic ranking performed by Qwen-VL-2B itself (self-labeling). The preference data curated through this process is then used to fine-tune Qwen-VL-2B. As presented in Table 3, we compare this aligned policy model against the original Qwen-VL-2B baseline and the same base model fine-tuned using a naïve RLAIF approach (which employs an external LLaVA-NeXT-34B model for ranking). The results clearly demonstrate the effectiveness and generalizability of applying TPR paradigm, even in a self-labeling setup, to improve performance on both hallucination and general benchmarks. This aligns with the conclusions outlined in our main results (see Section 4.2). Therefore, it is affirmed that the TPR paradigm is not confined to a specific model architecture like LLaVA. Its core data curation design can be effectively adapted to other models such as Qwen-VL-2B, leading to significant reductions in hallucinations while maintaining or improving general VLM capabilities.
Appendix E More Implementation Details
E.1 Data Curation
@@
Hyper-Parameters. For preference data curation, we utilize instruction prompt $x$ from RLHF-V [13] to guide the reference model $\pi_{\text{ref}}$ in generating $M=10$ candidate responses $\{y_{1},y_{2},\cdots,y_{M}\}$ for each image $I$. After these responses are decomposed into semantic units, $\pi_{\text{ref}}$ is queried with wh-questions derived from each unit to self-resample one additional candidate per topic. For the TPR-CL variant, 12,000 instances (60%) are constructed during “Warm-Up” stage, and the remaining 8,000 (40%) are constructed during “Hard-Mining” stage.
@@
Prompts. Within the TPR paradigm, the reference model $\pi_{\text{ref}}$ is prompted to perform several key operations: candidate response generation, response decomposition, topic clustering, in-context rewriting and the conversion of semantic units into both wh-questions and yes-no questions. The specific prompts utilized to guide $\pi_{\text{ref}}$ for these tasks are detailed in Table 4 and Table 5.
@@
Candidate Response Generation (Modified from RLHF-V)
We randomly select a following question to use as input for the reference model, prompting it to generate responses based on the image.
### Questions:
@@
Response Decomposition (Modified from RLAIF-V)
You are an expert in extracting facts from the given question-answer pair for an image. Your task is to extract and rewrite the facts mentioned in the answers into self-contained sentences. Exclude opinions or subjective statements.
You should present your result in the following format: ### Facts:
@@
Wh-Question Converting
You are an expert at modifying a given declarative sentence into a wh-question sentence. Your task is to modify the given declarative sentences one by one into a wh-question form. Do not change tenses or add extra content.
You should present your result in the following format: ### Converted questions:
@@
Yes-no Question Converting
You are an expert at modifying a given declarative sentence into a general question sentence. Your task is to modify the given declarative sentences one by one into a general question form. Do not change tenses or add extra content.
If the given declarative sentence contains not, no or negative meaning words, you need to check the modified general interrogative sentence to make sure that the generated general question sentence retains words with not, no or negative meaning words. You should present your result in the following format: ### Converted questions:
@@
Textual Consistency Evaluation
You are an expert at determining if the given two declarative sentences are consistent in textual semantics. Your task is to determine if the topic described in these two sentences are consistent. If you can confirm that two sentences are consistent, please output “consistent”. Otherwise, output “unrelated”.
### Declarative sentences:
@@
In-Context Rewriting
You are an expert at modifying a declarative answer with several tips. Your task is to modify the original answer, which is used to answer the question, based on the image and the provided tips. The given tips will relate to a specific part of the original answer, and you should use the tips to overwrite the corresponding part. If there is a conflict between the tips and the image, remember to follow the tips first.
You should make minimal modifications and maintain style and format with the original answer. Only output the modified answer. ### Tips:
@@ E.2 Evaluation
Appendix F Qualitative Case Studies
We provide qualitative case studies to illustrate two key aspects: (1) The differences between preference pairs generated by TPR and those generated by its curriculum learning variant, TPR-CL. See Figure 8 and 9. (2) The enhanced quality of responses from the policy model fine-tuned with TPR-curated data, particularly when compared against both the LLaVA-1.5-7B baseline and the larger LLaVA-NeXT-34B model. See Figure 10 and 11.
Appendix G Broader Impacts
TPR significantly enhances Vision Language Models by drastically reducing hallucinations, leading to more factually accurate and contextually coherent outputs, which in turn bolsters user trust. These improvements are pivotal for unlocking safer and more effective VLM applications across diverse domains, transforming these models into more reliable tools. On the other hand, a key consideration when applying TPR is ensuring that such targeted improvements in accuracy do not compromise the models’ broader general capabilities. Achieving this balance requires ongoing research and careful validation. Successfully addressing hallucinations is thus essential for the responsible development and widespread beneficial adoption of advanced VLM technology.
Appendix H Reproducibility
To ensure full reproducibility of our findings, we provide an anonymous code repository (“code.zip”) in supplementary materials. This repository includes all necessary components:
  • Scripts for setting up the environment, downloading the base models and the original data sources.
  • Scripts for preference data curation, including data source filtering and TPR-based curation.
  • Scripts for DPO training and subsequent evaluation.
Please refer to the “README.md” in “code.zip” for step-by-step instructions. We also provide an anonymous repository link as a backup.

### v4

@@ Appendix C Curriculum Learning
Effectiveness of Hard Mining. To investigate the impact of curriculum learning on model’s ability to avoid subtle errors, we conduct a quantitative analysis comparing effectiveness in reducing hallucination across different types of varying difficulties. The RefoMB benchmark [18] is used for this ablation, as its detailed categorizations allow us to pinpoint the benefits of curriculum learning more precisely. The results in Figure˜7 clearly demonstrate the progressive improvements achieved through our proposed TPR method and curriculum learning strategy. On challenging hallucinations for baseline LLaVA-1.5 model, i.e., “Quantities” (16.7%) and “Spatial Relations” (14.3%), the introduction of curriculum learning in TPR-CL yields further substantial improvements, providing an additional +20.8/+35.7 point increase over its greedy variant. For “Existence” and “Attributes”, where the greedy variant already performed strongly, TPR-CL still offers valuable refinements, with additional gains of +5.3 and +8.3 points respectively. These findings underscore the effectiveness of the hard mining component within our curriculum learning strategy. This is especially crucial for tackling more subtle and complex hallucinations related to quantities and spatial relationships, where the baseline model and even the greedy TPR approach show limitations.
@@
Appendix D Computational Cost
@@
table Computational Cost Breakdown. We provide the total computational cost of TPR generating full 20k dataset in each stage. vLLM is applied as inference engine in TPR for acceleration. All results are obtained using 8 NVIDIA A100 GPUs. Stage TPR TPR w/ vLLM Response Generation 10.27h 1.36h Decomposition 8.91h 3.57h Wh-question Convertion 7.95h 3.17h Self-resampling 7.48h 2.96h Topic Cluster 10.84h 6.71h Scoring & Ranking 23.43h 7.42h In-context Rewriting 2.17h 0.85h Total 71.05h 26.04h
[…]
To evaluate the practical viability of TPR, we conduct a comprehensive computational cost analysis. As illustrated in Figure˜7, TPR demonstrates the most favorable efficiency trajectory, achieving a rapid and substantial reduction in hallucinations at a minimal cost. This superior efficiency is rooted in our data generation pipeline, with a detailed breakdown provided in Figure˜8. The process of generating our 20k preference dataset on 8 NVIDIA A100 GPUs requires 71.05 hours, a time significantly reduced to just 26.04 hours with the integration of the vLLM [67] inference engine. This culminates in a highly efficient rate of 4.7 GPU-seconds per generated pair.
When compared with methods relying on proprietary APIs, TPR’s advantage is pronounced. As shown in Figure˜7, while a rewriting approach using GPT-4V yields a significant performance improvement, it comes at a much higher price. TPR achieves a comparable level of hallucination reduction at approximately one-third of the estimated API cost required by the GPT-4V approach (roughly \$10 for TPR vs. over \$30 for GPT-4V). Furthermore, TPR is also more cost-effective than other open-source alternatives like RLAIF-V. Under the same hardware conditions, generating a 22k dataset with the official RLAIF-V implementation takes approximately 66 hours. Although RLAIF-V involves fewer distinct stages, its process is dominated by a computationally expensive scoring phase using a 34B parameter labeler model. More critically, its iterative refinement strategy introduces significant overhead: the policy model must be repeatedly retrained between data generation rounds, a resource-intensive process that cannot be accelerated by inference engines like vLLM. This inherent inefficiency explains the gentler decline in its performance-to-cost curve shown in Figure˜7.
In summary, these findings clearly indicate that TPR not only achieves state-of-the-art hallucination reduction but does so with significantly greater computational and financial efficiency than existing methods, highlighting its practical value for large-scale VLM alignment.
Appendix E Qualitative Case Studies
We provide qualitative case studies to illustrate two key aspects: (1) The differences between preference pairs generated by TPR and those generated by its curriculum learning variant, TPR-CL. See Figure˜9 and 10. (2) The enhanced quality of responses from the policy model fine-tuned with TPR-curated data, particularly when compared against both the LLaVA-1.5-7B baseline and the larger LLaVA-NeXT-34B model. See Figure˜11 and 12.
Appendix F More Implementation Details
F.1 Data Curation
@@
Hyper-Parameters. For preference data curation, we utilize instruction prompt $x$ from RLHF-V [13] to guide the reference model $\pi_{\text{ref}}$ in generating $M=10$ candidate responses $\{y_{1},y_{2},\cdots,y_{M}\}$ for each image $I$. After these responses are decomposed into semantic units, $\pi_{\text{ref}}$ is queried with wh-questions derived from each unit to self-resample one additional candidate per topic. For the TPR-CL variant, 12,000 instances (60%) are constructed during “Warm-Up” stage, and the remaining 8,000 (40%) are constructed during “Hard-Mining” stage. For initial candidate responses sampling and intra-topic resampling, we follow RLHF-V [13] to set the parameters as: temperature=0.7,top_p=0.95,do_sample=True. Other parameters like top_k are left at their default values.
@@
Prompts. Within the TPR paradigm, the reference model $\pi_{\text{ref}}$ is prompted to perform several key operations: candidate response generation, response decomposition, topic clustering, in-context rewriting and the conversion of semantic units into both wh-questions and yes-no questions. The specific prompts utilized to guide $\pi_{\text{ref}}$ for these tasks are detailed in Table˜3 and Table˜4.
@@
Candidate Response Generation (Modified from RLHF-V) We randomly select a following question to use as input for the reference model, prompting it to generate responses based on the image. ### Questions:
@@
Response Decomposition (Modified from RLAIF-V) You are an expert in extracting facts from the given question-answer pair for an image. Your task is to extract and rewrite the facts mentioned in the answers into self-contained sentences. Exclude opinions or subjective statements. You should present your result in the following format: ### Facts:
@@
Wh-Question Converting You are an expert at modifying a given declarative sentence into a wh-question sentence. Your task is to modify the given declarative sentences one by one into a wh-question form. Do not change tenses or add extra content. You should present your result in the following format: ### Converted questions:
@@
Yes-no Question Converting You are an expert at modifying a given declarative sentence into a general question sentence. Your task is to modify the given declarative sentences one by one into a general question form. Do not change tenses or add extra content. If the given declarative sentence contains not, no or negative meaning words, you need to check the modified general interrogative sentence to make sure that the generated general question sentence retains words with not, no or negative meaning words. You should present your result in the following format: ### Converted questions:
@@
Textual Consistency Evaluation You are an expert at determining if the given two declarative sentences are consistent in textual semantics. Your task is to determine if the topic described in these two sentences are consistent. If you can confirm that two sentences are consistent, please output “consistent”. Otherwise, output “unrelated”. ### Declarative sentences:
@@
In-Context Rewriting You are an expert at modifying a declarative answer with several tips. Your task is to modify the original answer, which is used to answer the question, based on the image and the provided tips. The given tips will relate to a specific part of the original answer, and you should use the tips to overwrite the corresponding part. If there is a conflict between the tips and the image, remember to follow the tips first. You should make minimal modifications and maintain style and format with the original answer. Only output the modified answer. ### Tips:
@@ F.2 Evaluation
Appendix G More Ablation Studies
Larger Base Model. To investigate the scalability of our method, we applied TPR-CL to both 7B and 13B variants of the base model across data scales from 2k to 20k. As illustrated in Figure˜8, the 13B model consistently outperforms its 7B counterpart at every data point, achieving better hallucination scores in both ObjHal [55] and AMBER [56] benchmarks. This result confirms that the performance gains from TPR are complementary to the inherent capabilities of larger models, highlighting its strong scalability.
Self-Labeling: Using the Reference Model Itself as Labeler. The approach of scoring fine-grained semantic units within our intra-topic ranking mechanism potentially alleviates the demand for an exceptionally capable labeler model [18]. This prompted an exploration into the effectiveness of using the reference model itself for scoring and ranking after self-resampling, thereby investigating the boundaries of model self-improvement. Accordingly, we introduce TPR-SL (Self-Labeling), a variant of TPR that leverages the base LLaVA-1.5-7B model [7] as its own preference labeler, instead of a more powerful one like LLaVA-NeXT-34B [4] used in our main experiments. The performance of TPR-SL compared to the LLaVA-1.5-7B baseline is presented in Table˜5. Specifically, TPR-SL achieves substantial improvements over the baseline LLaVA-1.5-7B across both hallucination mitigation and general VLM capabilities. These findings are significant as they highlight the effectiveness of TPR in a self-labeling scenario. It validates a practical and efficient paradigm for VLM self-improvement that minimizes reliance on human annotations or powerful external models.
Hallucination Benchmarks General Benchmarks
Model ObjHal MMHal AMBER RefoMB POPE LLaVA-B MMstar
$_{\text{s}} \downarrow CH$$_{\text{i}} \downarrow$ Score $\uparrow$ Hall. $\downarrow$ Acc. $\uparrow$ F1 $\uparrow$ Trust. $\uparrow$ Win. $\uparrow$ F1 $\uparrow$ Overall $\uparrow$ Overall $\uparrow$
LLaVA-1.5-7B 53.6 25.2 2.36 51.0 73.5 77.6 30.8 12.1 85.9 59.7 30.3
+ TPR-SL 5.8 3.0 2.67 44.8 81.7 86.7 53.5 30.3 86.1 73.7 32.8
Qwen-VL-2B [5] 42.4 36.3 2.85 47.9 68.3 81.0 48.5 20.7 86.5 83.2 29.1
+ Naive RLAIF 36.5 31.7 2.81 49.0 70.9 81.8 52.5 23.7 87.0 83.2 30.3
+ TPR-SL-2B 19.5 13.7 2.98 43.8 74.3 84.3 56.6 28.3 87.1 83.8 31.2
Generalization on Different Model Architecture. To verify the generalizability of the TPR paradigm, we apply it to a different model architecture, Qwen-VL-2B [5]. Specifically, Qwen-VL-2B serves as the reference model for both generating topic-level alternatives and subsequent selective replacement, with intra-topic ranking performed by Qwen-VL-2B itself (self-labeling). The preference data curated through this process is then used to fine-tune Qwen-VL-2B. As presented in Table˜5, we compare this aligned policy model against the original Qwen-VL-2B baseline and the same base model fine-tuned using a naïve RLAIF approach (which employs an external LLaVA-NeXT-34B model for ranking). The results clearly demonstrate the effectiveness and generalizability of applying TPR paradigm, even in a self-labeling setup, to improve performance on both hallucination and general benchmarks. This aligns with the conclusions outlined in our main results (see Section 4.2). Therefore, it is affirmed that the TPR paradigm is not confined to a specific model architecture like LLaVA. Its core data curation design can be effectively adapted to other models such as Qwen-VL-2B, leading to significant reductions in hallucinations while maintaining or improving general VLM capabilities.
Appendix H Broader Impacts
The research presented in this paper aims to enhance the reliability of Vision Language Models (VLMs). By significantly reducing visual hallucinations, TPR leads to more factually accurate and contextually coherent outputs, which in turn can bolster user trust. These improvements are pivotal for unlocking safer and more effective VLM applications across diverse domains, from assistive technologies to educational tools, transforming these models into more reliable instruments. However, the deployment of this technology require careful consideration of potential societal impacts and technical trade-offs.
Potential for Misuse. A primary concern is that the synthetic “rejected” responses ($y_{l}$) generated by TPR, which are designed to contain plausible hallucinations, could be isolated and misused for creating disinformation if taken out of context. To mitigate this risk, we are committed to responsible asset release. Any public distribution of our curated preference dataset will be governed by a strict, research-focused license prohibiting malicious use. Furthermore, the dataset will be accompanied by comprehensive documentation that explicitly warns users that the negative examples are synthetic, generated solely for model alignment, and are factually incorrect and unsuitable for any other application.
Environmental and Computational Cost. Large-scale data curation and model alignment inevitably have an environmental cost due to the required computational resources. As detailed in our analysis (see Appendix˜D), TPR is applied during the post-training alignment phase, making its environmental impact substantially lower than that of pre-training a VLM from scratch. Crucially, our work emphasizes data efficiency. By generating high-quality preference pairs that provide strong learning signals, TPR achieves state-of-the-art performance with less data than many competing methods, thereby promoting a smaller environmental footprint for VLM alignment.
Technical Safeguards. Besides, a key technical consideration when applying TPR is ensuring that targeted improvements in factuality do not compromise the model’s broader general capabilities. Although our experiments (see Table˜1) show that our approach maintains or even enhances general performance, achieving this balance requires ongoing research and careful validation.
Successfully navigating these challenges is essential for the responsible development and widespread beneficial adoption of advanced VLM technology.